OpenStudy (anonymous):

okay, so how do you factor 64x^3-8?

OpenStudy (anonymous):

64=4^3 8=2^3 use\[x^3-y^3=(x-y)(x^2+xy+y^2)\]

OpenStudy (anonymous):

that doesn't exactly help me..

OpenStudy (saifoo.khan):

There's a formula for solving problems like these. @jonask wrote that formula above.

OpenStudy (anonymous):

So for example \[1000x^3 - 125y^3\to (10x)^3-(5y)^3\] Using SOAP \[(10x-5y)((10x)^2+(10x)(5y)+(5y)^2)\] and then you clean up thingsl like 10x5y=50xy and stuff like that

OpenStudy (anonymous):

okay, the answer is 8(2x-1)(4x^2+2x+1). I'm not sure how to get that...

OpenStudy (anonymous):

Were you able to apply the SOAP method to your problem? could you convert \[64x^3\rightarrow ()^3\] ?

OpenStudy (anonymous):

yea, I got 64x^3-8=(4x)^3-2^3=(4x-2)(16x^2+8x+4). I don't know what to do from there

OpenStudy (anonymous):

factor factor factor. 1) what # can you factor from (4x-2) ? 2) what # can you factor from (16x^2+8x+4)?

OpenStudy (anonymous):

BTW, you did the SOAP correctly:)

OpenStudy (anonymous):

err....I'm not totally sure...

OpenStudy (anonymous):

I'll do the first one and you can do the second. (4x+2) = 2(2x+1) since you can divide 4 and 2 by 2:)

OpenStudy (anonymous):

What is the larges integer number that divides 16, 8, and 4?

OpenStudy (anonymous):

4, and then you multiply two and 4 and that's where the 8 comes from! I see!

OpenStudy (anonymous):

Yay! Remember, next to perseverence (sp?), Factoring is probably the most important skill for math

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

use polynomial division procedure:D